OpenStudy (doggiebounce):

What is the average kinetic energy of 1 mole of a gas at 125 Kelvin? (R = 8.314 J/K-mol) A. 1.04 x 10^3 J B. 3.12 x 10^3 J C. 1.56 x 10^3 J D. 1.77 x 10^3 J

OpenStudy (chmvijay):

do you know equation for average kinetic energy

sam (.sam.):

From kinetic theory, \[E_K=\frac{1}{2}m<c^2>=\frac{3}{2}k_BT\] --------------------------------------- We can say that \[E_K=\frac{3}{2}k_BT\] \(k_B=Boltzmann~Constant=1.38 \times 10^{-23}m^2kgs^{-2}K^{-1}\) Substitute in \[E_K=\frac{3}{2}(1.38 \times 10^{-23})(125)=?\] Multiply that by avogadro number \(N_A\) gives average kinetic energy

sam (.sam.):

@chmvijay do you have a different approach? I can't think of one relating R

OpenStudy (chmvijay):

ur right i is 3KT/2

sam (.sam.):

yeah but the problem is the question includes (R = 8.314 J/K-mol)

OpenStudy (anonymous):

Did anyone ever figure out how to do this?